# ALPCAHUS: Subspace Clustering for Heteroscedastic Data

### v2

@@ III-A3 Subspace Clustering via Thresholding (TSC)
In the example to follow, assume that there exists two unique 1-dimensional subspaces. Given two data samples $\bm{y}_{i}$ and $\bm{y}_{j}$, intuitively their dot product $\langle\bm{y}_{i},\bm{y}_{j}\rangle$ will be high if $c_{i}=c_{j}$ meaning they belong to the same subspace. Likewise, if $c_{i}\neq c_{j}$, then $\langle\bm{y}_{i},\bm{y}_{j}\rangle=0$ assuming orthogonal subspaces. In non-orthogonal scenarios, $\langle\bm{y}_{i},\bm{y}_{j}\rangle$ will be relatively small under $c_{i}\neq c_{j}$ compared to $c_{i}=c_{j}$. Using this idea, in more general D-dimensional subspace settings, TSC constructs a matrix $\bm{Z}\in\mathbb{R}^{N\times N}$ such that
@@
where $\|\bm{W}_{i,:}\|_{0}=q$ is the $l_{0}$ pseudo-norm. In other words, one constructs $\bm{W}_{i,:}$ using the q nearest neighbors that correspond to the highest magnitude values. Then, spectral clustering is applied to $\bm{W}$ to find the subspace clustering associations.
III-A4 Ensemble $K$-Subspaces (EKSS)
Iterative methods are based on the idea of alternating between cluster assignments and subspace basis approximation, with KSS being highly prominent [25]. The KSS algorithm seeks to solve the following optimization problem:
$\operatorname*{arg\,min}_{\mathcal{C},\,\mathcal{U}}\sum_{k=1}^{K}\sum_{c_{i}=k,\forall i}\|\bm{y}_{i}-\bm{U}_{k}\bm{U}_{k}^{T}\bm{y}_{i}\|_{2}^{2}.$ (11)
The goal is to minimize the sum of residual norms by alternating between performing PCA on each cluster to update $\mathcal{U}$ and using the subspace bases to calculate new cluster assignments $\mathcal{C}$ in a similar fashion to the $K$-means algorithm.
The quality of the solution depends highly on the initialization. Recent work provides convergence guarantees and spectral initialization schemes that provably perform better than random initialization [26]. However, this problem (11), as shown in [27], is NP-hard. Further, it is prone to local minima [28]. To overcome this, consensus clustering [29] is a tool that leverages information from many trials and combines results together. This approach, known as Ensemble KSS (EKSS) [16] in this context, creates an affinity matrix whose $(i,j)$th entry represents the number of times the two points were clustered together in a trial. Then, spectral clustering is performed on the affinity matrix to get the final clustering from the many base clusterings. The use of PCA makes it challenging to learn clusters and subspace bases in the heteroscedastic regime since PCA implicitly assumes the same noise variance across all samples [30]. The proposed ALPCAHUS approach in Sec. IV builds on KSS, and its ensemble version, by generalizing (11) to the heteroscedastic regime.
III-B Single Subspace Heteroscedastic PCA Method
Before extending (11) to the heteroscedastic setting, a review of how LR-ALPCAH [17] solves for a single subspace ($K=1$) in the heteroscedastic setting using the model described in (1) is necessary. For the measurement model $\bm{y}_{i}\sim\mathcal{N}(\bm{x}_{i},\nu_{i}\bm{I})$ in (1), the probability density function for a single data sample $\bm{y}_{i}$ is
After further manipulation, the negative log-likelihood is expressed in matrix notation as
$\frac{1}{2}\|(\bm{Y}-\bm{X})\bm{\Pi}^{-1/2}\|_{\mathrm{F}}^{2}+\frac{D}{2}\log|\bm{\Pi}|$ (13)
where $\bm{\Pi}=\text{diag}(\nu_{1},\ldots,\nu_{N})$. To reduce computation and enforce a low-rank solution, LR-ALPCAH [17] took inspiration from the matrix factorization literature [31] and factorized $\bm{X}\in\mathbb{R}^{D\times N}\approx\bm{L}\bm{R}^{\prime}$ where $\bm{L}\in\mathbb{R}^{D\times\hat{d}}$ and $\bm{R}\in\mathbb{R}^{N\times\hat{d}}$ for some rank estimate $\hat{d}$. Using this idea, LR-ALPCAH estimates $\bm{X}$ by solving for $\bm{L}$ and $\bm{R}$ in the following optimization problem
The next section, Sec. IV, combines ideas from (11) and (14) to tackle the general union of subspaces setting ($K>1$).
III-C Other heteroscedastic models
This paper focuses on heteroscedastic noise across the data samples. There are other subspace learning methods in the literature that explore heteroscedasticity in different ways. For example, HeteroPCA considers heteroscedasticity across the feature space [32]. One possible application of that model is for data that consists of sensor information with multiple devices that naturally have different levels of precision and signal to noise ratio (SNR). Another heterogeneity model considers the noise to be homoscedastic and instead assumes that the signal itself is heteroscedastic [33]. That work considers clustering applications where the power fluctuating (i.e., heteroscedastic) signals are embedded in white Gaussian noise. We exclude comparisons with those methods since their models are very different. These different models each have their own applications.
IV Proposed Subspace Clustering Method
For notational simplicity, let $\bm{Y}_{k}\triangleq\bm{Y}_{\mathcal{C}_{k}}\in\mathbb{R}^{D\times N_{k}}$ denote the submatrix of $\bm{Y}$ having columns corresponding to data samples that are estimated to belong in the $k$th subspaces, i.e., $\bm{Y}_{k}\triangleq\bm{Y}_{\mathcal{C}_{k}}\triangleq\text{matrix}(\{\bm{y}_{i}\,:\,c_{i}=k\})$. This notation applies similarly to other matrices such as $\bm{\Pi}_{k}\triangleq\bm{\Pi}_{\mathcal{C}_{k}}\triangleq\text{diag}(\{\nu_{i}\,:\,c_{i}=k\})$. For the union of subspace measurement model (2), we generalize LR-ALPCAH (14) with the following optimization problem
$\operatorname*{arg\,min}_{\mathcal{L},\,\mathcal{R},\,\bm{\Pi},\,\mathcal{C}}\sum_{k=1}^{K}$ $\frac{1}{2}\left\|(\bm{Y}_{k}-\bm{L}_{k}\bm{R}^{T}_{k})\bm{\Pi}^{-1/2}_{k}\right\|_{\mathrm{F}}^{2}+\frac{D}{2}\log|\bm{\Pi}_{k}|$ (15)
where $\mathcal{C},\bm{\Pi},\mathcal{L},\mathcal{R}$ denote the sets of estimated clusters, noise variances, and factorized matrices respectively for each cluster $k=1,\ldots,K$. Specifically $\mathcal{L}\triangleq\{\bm{L}_{1},\ldots,\bm{L}_{K}\}$ and $\mathcal{R}\triangleq\{\bm{R}_{1},\ldots,\bm{R}_{K}\}$. Our algorithm for solving (15) is called ALPCAHUS (ALPCAH for Union of Subspaces). Similar to KSS, ALPCAHUS alternates between updating subspace bases given cluster estimates and updating cluster estimates by projection given subspace bases estimates. We solve (15) with the alternating minimization technique. The method alternates between basis estimation, $\forall k\in\{1,\ldots,K\}$ solve
$\operatorname*{arg\,min}_{\bm{L}_{k},\,\bm{R}_{k},\,\bm{\Pi}_{k}}$ $\frac{1}{2}\left\|(\bm{Y}_{k}-\bm{L}_{k}\bm{R}_{k}^{T})\bm{\Pi}_{k}^{-1/2}\right\|_{\mathrm{F}}^{2}$
@@ IV Proposed Subspace Clustering Method
and data sample reassignment, $\forall i\in\{1,\ldots,N\}$ solve
@@
$c_{i}^{(t_{2}+1)}\leftarrow c_{i}^{(t_{2})}\text{ if }\exists\tilde{k}\in\mathcal{S}_{J_{i}}\neq c_{i}^{(t_{2})}\text{ such that }$
$J_{i}(c_{i}^{(t_{2})})=J_{i}(\tilde{k})\text{ occurs}.$ (17)
The solution to (16) is described in [17] with convergence theory. For completeness and notational consistency, the updates are included in (18) (19) (21). For the $(t_{1}+1)$ iteration, given current $t_{1}$ iterates, the updates are given as
$\bm{L}^{(t_{1}+1)}_{k}$ $=\operatorname*{arg\,min}_{\bm{L}_{k}}f(\bm{L}_{k},\bm{R}^{(t_{1})}_{k},\bm{\Pi}^{(t_{1})}_{k})$
$=\bm{Y}_{k}\bm{\Pi}^{(t_{1})}_{k}\bm{R}^{(t_{1})}_{k}(\bm{R}^{(t_{1})^{T}}_{k}\bm{\Pi}^{(t_{1})}_{k}\bm{R}^{(t_{1})}_{k})^{-1}$ (18)
$\bm{R}^{(t_{1}+1)}_{k}$ $=\operatorname*{arg\,min}_{\bm{R}_{k}}f(\bm{L}^{(t_{1})}_{k},\bm{R}_{k},\bm{\Pi}^{(t_{1})}_{k})$
$=\bm{Y}^{T}_{k}\bm{L}^{(t_{1})}_{k}(\bm{L}^{(t_{1})^{T}}_{k}\bm{L}^{(t_{1})}_{k})^{-1}.$ (19)
Let the following function be defined
$p_{k}(j)\triangleq\max(\alpha,|\mathcal{C}_{k}|^{-1}\|(\bm{Y}_{k}-\bm{L}^{(t_{1})}_{k}\bm{R}^{(t_{1})^{T}}_{k})\bm{e}_{j}\|_{2}^{2})$ (20)
where $\bm{e}_{j}$ denotes the $j$th standard canonical basis vector that is used to select the $j$th column of some matrix and $\alpha>0$ is some noise variance threshold parameter. Then, the noise variance matrix update is given as
$\bm{\Pi}^{(t_{1}+1)}_{k}$ $=\operatorname*{arg\,min}_{\bm{\Pi}_{k}}f(\bm{L}^{(t_{1})}_{k},\bm{R}^{(t_{1})}_{k},\bm{\Pi}_{k})\implies$
$=\text{diag}(p_{k}(1),\ldots,p_{k}(|\mathcal{C}_{k}|)).$ (21)
Because $\bm{\Pi}_{k}$ is a diagonal matrix, the $\bm{\Pi}_{k}\succeq\alpha\bm{I}$ majorization condition in (16) being equivalent to $\forall i,\lambda_{i}(\bm{\Pi}_{k})\geq\alpha$ implies that $\forall i,\nu_{i}>\alpha$. This leads to a projection to the positive orthant, i.e., the $\max(\alpha,\cdot)$ condition in (20). This condition is necessary to ensure convergence as proven in [17, Thm. 2] and used in this work to further prove ALPCAHUS convergence in Thm. 1 using some $\alpha\in\mathbb{R}_{+}$.
@@
The solution to (17) involves the following procedure. Given a data sample $\bm{y}_{i}$, find the lowest subspace projection residual out of all subspaces and assign it to $c_{i}^{(t_{2}+1)}$ at the $(t_{2}+1)$ iteration. In the event of a tie, meaning there is more than one subspace that has equal residual, copy the past label $c_{i}^{(t_{2})}$ to $c_{i}^{(t_{2}+1)}$. In (17), $\mathcal{S}_{J_{i}}$ refers to the set of possible $k$ that achieves the lowest cost $J_{i}(k)$, i.e., $J_{i}(k_{1})=J_{i}(k_{2})\text{ for }k_{1}\neq k_{2}$. By doing this, cycling between cluster labels is prevented for all data samples. This cluster reassignment criteria will play a big role in ensuring convergence as proven in Thm. 1.
@@
To further improve clustering performance, consensus clustering can be leveraged over many trials. Initially, we tried using this approach with only one trial as seen in Fig. 2(b) and ALPCAHUS ($B=1$) result in Fig. 6(a). However, since $K$-subspaces in general is sensitive to initialization, during experimentation there was great success in using a consensus approach with more than one trial as shown in Fig. 2(d) and ALPCAHUS ($B=16$) result in Fig. 6(a).
Algorithm ALPCAHUS (unknown noise variances, unknown noise grouping)
  Input: $\bm{Y}\in\mathbb{R}^{D\times N}$: data, $K\in\mathbb{Z}^{+}$: number of subspaces, $\{\hat{d}_{k}\in\mathbb{Z}^{+}\hskip 5.69054pt\forall k\in\{1,\ldots,K\}\}$: candidate dimension for all clusters, $q\in\mathbb{Z}^{+}$: threshold parameter, $B\in\mathbb{Z}^{+}$: base clusterings, $T_{1}\in\mathbb{Z}^{+}$: LR-ALPCAH iterations, $T_{2}\in\mathbb{Z}^{+}$: maximum alternating updates (cluster reassignments)
  Output: $\mathcal{C}=\{c_{1},\dots,c_{K}\}$: clusters of $\bm{Y}$
  for $b=1,\dots,B$ (in parallel)  do
     $\mathcal{C}_{k}\sim\{1,\ldots,N\}\hskip 5.69054pt\text{s.t.}\hskip 5.69054pt|\mathcal{C}_{k}|\approx\frac{N}{\bar{K}}$ for $k=1,\dots,K$ Initialize clusters randomly without replacement
     for $k=1,\dots,K$ (in parallel)  do
        Apply LR-ALPCAH from [17] with rank $\hat{d}_{k}$ matrices for $\bm{L}_{k}$ and $\bm{R}_{k}$
        $\bm{L}_{k},\bm{R}_{k},\bm{\Pi}_{k}\leftarrow$ Initialize $\bm{\Pi}_{k}=\bm{I}$ and compute (18) (19) (21) using $\bm{Y}_{k}$ for $T_{1}$ iterations
     end for
     for $t_{2}=1,\dots,T_{2}$ (in sequence)  do
        for $k=1,\dots,K$ (in parallel)  do
           $\bm{L}_{k},\bm{R}_{k},\bm{\Pi}_{k}\leftarrow$ Compute (18) (19) (21) using $\bm{Y}_{k}$ for $T_{1}$ iterations Apply LR-ALPCAH from [17] with rank $\hat{d}_{k}$
        end for
        $\mathcal{C}_{k}\leftarrow\{\forall\bm{y}_{i}\in\bm{Y}:$ Solve (17) with $\bm{U}_{k}\text{ extracted from SVD}(\bm{L}_{k}),\forall k\}$ Update current cluster estimates by projection
        Compute cost function value in (31) and stop if $\text{cost}^{(t_{1}+1,t_{2}+1)}=\text{cost}^{(t_{1},t_{2})}$ Stopping criteria
     end for
     $\mathcal{C}^{(b)}\leftarrow\mathcal{C}_{k}=\{c_{1},\dots,c_{N}\}$ Collect results from all trials
  end for
  $\bm{W}_{i,j}\leftarrow$ Form $\bm{W}$ using (22) Form affinity matrix of similar clusterings
  for $i=1,\dots,N$ (in parallel)  do
     $\bm{Z}^{\text{row}}\leftarrow$ Threshold rows by solving (23) Retain top $q$ elements for each row
     $\bm{Z}^{\text{col}}\ \leftarrow$ Threshold columns by solving (24) Retain top $q$ elements for each column
  end for
  $\bm{W}\leftarrow\frac{1}{2}\left(\bm{Z}^{\text{row}}+\bm{Z}^{\text{col}}\right)$ Average affinity matrix
  $\mathcal{C}\leftarrow$ Perform spectral clustering via (6) and apply $K$-means on the rows of $\hat{H}$ Final clustering

### v3

@@ III-A3 Subspace Clustering via Thresholding (TSC)
In the example to follow, assume that there are two distinct 1-dimensional subspaces. Given two data samples $\bm{y}_{i}$ and $\bm{y}_{j}$, intuitively their dot product $\langle\bm{y}_{i},\bm{y}_{j}\rangle$ will be high if $c_{i}=c_{j}$, meaning they belong to the same subspace. Likewise, if $c_{i}\neq c_{j}$, then $\langle\bm{y}_{i},\bm{y}_{j}\rangle=0$ if the subspaces are orthogonal. In non-orthogonal scenarios, one expects $|\langle\bm{y}_{i},\bm{y}_{j}\rangle|$ to be smaller when $c_{i}\neq c_{j}$ than when $c_{i}=c_{j}$. Using this idea, in more general D-dimensional subspace settings, TSC constructs a matrix $\bm{Z}\in\mathbb{R}^{N\times N}$ that describes similarity such that
$\bm{Z}_{ij}=\exp(-2\arccos(|\langle\bm{y}_{i},\bm{y}_{j}\rangle|))\hskip 5.69054pt\text{s.t.}\hskip 5.69054pt\hskip 5.69054pt\bm{Z}_{ii}=0\hskip 2.84526pt\forall i.$ (10)
$\bm{W}=\operatorname*{arg\,min}_{\bm{W}}\left\|\bm{W}-\bm{Z}\right\|_{\mathrm{F}}^{2}\hskip 5.69054pt\text{s.t.}\hskip 5.69054pt\|\bm{W}_{i,:}\|_{0}=q\hskip 5.69054pt\forall i$ (11)
where $\|\bm{W}_{i,:}\|_{0}=q$ is the $l_{0}$ pseudo-norm that counts the number of nonzero entries. In other words, one constructs $\bm{W}_{i,:}$ using the q nearest neighbors that correspond to the highest magnitude values. Then, spectral clustering is applied to $\bm{W}$ to find the cluster associations.
III-A4 Ensemble $K$-Subspaces (EKSS)
Iterative methods are based on the idea of alternating between cluster assignments and subspace basis approximation, with KSS being highly prominent [kplane]. The KSS algorithm seeks to solve
$\operatorname*{arg\,min}_{\mathcal{C},\,\mathcal{U}}\sum_{k=1}^{K}\sum_{c_{i}=k,\forall i}\|\bm{y}_{i}-\bm{U}_{k}\bm{U}_{k}^{\prime}\bm{y}_{i}\|_{2}^{2}.$ (12)
The goal is to minimize the sum of residual norms by alternating between performing PCA on each cluster to update $\mathcal{U}$ and using the subspace bases to calculate new cluster assignments $\mathcal{C}$ in a similar fashion to the $K$-means algorithm. The quality of the solution depends highly on the initialization. Recent work provides convergence guarantees and spectral initialization schemes that provably perform better than random initialization [kss_convergence]. However, this problem (12), is known to be NP-hard [gitlin2018improving]. Further, it is prone to local minima [timmerman2013subspace]. To overcome this, consensus clustering [consensus_clustering] is a tool that leverages information from many trials and combines results together. This approach, known as Ensemble KSS (EKSS) [ekss], creates an affinity matrix whose $(i,j)$th entry represents the number of times the two points were clustered together in a trial. Then, spectral clustering is performed on the affinity matrix to get the final clustering from the many base clusterings. The use of PCA makes it challenging to learn clusters and subspace bases in the heteroscedastic regime since PCA implicitly assumes the same noise variance across all samples [heppcat]. The proposed ALPCAHUS approach in Sec. IV builds on KSS, and its ensemble version, by generalizing (12) to the heteroscedastic regime while simultaneously learning data quality.
III-B Other heteroscedastic models
This paper focuses on heteroscedastic noise across the data samples. There are other subspace learning methods in the literature that explore heteroscedasticity in different ways. For example, HeteroPCA considers heteroscedasticity across the feature space [hetero_pca]. One possible application of that model is for data that consists of sensor information with multiple devices that naturally have different levels of precision and signal to noise ratio. Another heterogeneity model considers the noise to be homoscedastic and instead assumes that the signal itself is heteroscedastic [collas2021probabilistic]. That work considers applications where the power fluctuating signals, i.e., heteroscedastic signals, are embedded in white Gaussian noise. However, to the authors’ knowledge, there are no algorithms for subspace clustering that consider feature-space heteroscedasticity. We exclude comparisons with these methods since their models are very different and each have their own applications.
III-C Single Subspace Heteroscedastic PCA
Before extending (12) to the heteroscedastic setting, we will review how LR-ALPCAH [alpcah_journal] solves for a single subspace ($K=1$) in the heteroscedastic setting using the model described in (1). For the measurement model $\bm{y}_{i}\sim\mathcal{N}(\bm{x}_{i},\nu_{i}\bm{I})$ in (1), the probability density function for a single data sample $\bm{y}_{i}$ following a Gaussian distribution is easily expressed as
For independent samples, after dropping constants, the joint log likelihood of all data $\{\bm{y}_{i}\}_{i=1}^{N}$ is the following
$\sum_{i=1}^{N}-\frac{1}{2}\log|\nu_{i}\bm{I}|-\frac{1}{2}(\bm{y}_{i}-\bm{x}_{i})^{\prime}(\nu_{i}\bm{I})^{-1}(\bm{y}_{i}-\bm{x}_{i}).$ (14)
Let $\bm{\Pi}=\mathrm{diag}(\nu_{1},\ldots,\nu_{N})\in\mathbb{R}^{N\times N}$ be a diagonal matrix representing the typically unknown noise variances. Then, the negative log likelihood in matrix form is
$\frac{D}{2}\log|\bm{\Pi}|+\frac{1}{2}\text{Tr}[(\bm{Y}-\bm{X})\bm{\Pi}^{-1}(\bm{Y}-\bm{X})^{\prime}].$ (15)
After further manipulation by trace lemmas, we rewrite the negative log-likelihood as
$\frac{1}{2}\|(\bm{Y}-\bm{X})\bm{\Pi}^{-1/2}\|_{\mathrm{F}}^{2}+\frac{D}{2}\log|\bm{\Pi}|.$ (16)
Our previous work used a functional operator similar to the nuclear norm to regularize $\bm{X}$ and encourage a low-rank solution by penalizing the tail sum of singular values [locallylowrank] [alpcah]. However, this method named ALPCAH was relatively slow due to SVD operations every iteration. To reduce computation and enforce a low-rank solution, the LR-ALPCAH variant [alpcah_journal] took inspiration from the matrix factorization literature [matrix_factorization] and factorized $\bm{X}\in\mathbb{R}^{D\times N}\approx\bm{L}\bm{R}^{\prime}$ where $\bm{L}\in\mathbb{R}^{D\times\hat{d}}$ and $\bm{R}\in\mathbb{R}^{N\times\hat{d}}$ for some rank estimate $\hat{d}$. Using this idea, LR-ALPCAH estimates $\bm{X}$ by solving for $\bm{L}$ and $\bm{R}$, jointly with noise variance matrix $\bm{\Pi}$, by the following
This comes from a modified model of (1) described by
$\bm{y}_{i}=\bm{L}\bm{r}_{i}+\bm{\epsilon}_{i},\quad\bm{\epsilon}_{i}\sim\mathcal{N}(\bm{0},\nu_{i}\bm{I})$ (18)
where $\bm{r}_{i}$ denotes the $i$th column of $\bm{R}$. Sec. IV combines ideas from (12) and (17) to tackle the union of subspaces setting.
IV Proposed Subspace Clustering Method
For notational simplicity, let $\bm{Y}_{k}=\bm{Y}_{\mathcal{C}_{k}}\in\mathbb{R}^{D\times N_{k}}$ denote the submatrix of $\bm{Y}$ having columns corresponding to data samples that are estimated to belong in the $k$th subspace, i.e., $\bm{Y}_{k}=\bm{Y}_{\mathcal{C}_{k}}=\text{matrix}(\{\bm{y}_{i}\,:\,c_{i}=k\})$, where $\mathcal{C}_{k}=\{i\,:\,c_{i}=k\}$ must be determined. This notation applies similarly to other matrices such as $\bm{\Pi}_{k}=\bm{\Pi}_{\mathcal{C}_{k}}=\text{diag}(\{\nu_{i}\,:\,c_{i}=k\})$. For the union of subspace measurement model in (2), we generalize LR-ALPCAH (17) by proposing the following cost function
$\operatorname*{arg\,min}_{\mathcal{L},\,\mathcal{R},\,\bm{\Pi},\,\mathcal{C}}\underbrace{\sum_{k=1}^{K}\overbrace{\frac{1}{2}\left\|(\bm{Y}_{k}-\bm{L}_{k}\bm{R}_{k}^{\prime})\bm{\Pi}^{-1/2}_{k}\right\|_{\mathrm{F}}^{2}+\frac{D}{2}\log|\bm{\Pi}_{k}|}^{f_{k}(\bm{L}_{k},\bm{R}_{k},\bm{\Pi}_{k};\bm{Y}_{k})}}_{f(\mathcal{L},\mathcal{R},\bm{\Pi},\mathcal{C})}$ (19)
where $\mathcal{C},\bm{\Pi},\mathcal{L},\mathcal{R}$ denote the sets of estimated clusters, noise variances, and factorized matrices respectively for each cluster $k=1,\ldots,K$. Specifically, $\mathcal{L}\triangleq\{\bm{L}_{1},\ldots,\bm{L}_{K}\}$ and $\mathcal{R}\triangleq\{\bm{R}_{1},\ldots,\bm{R}_{K}\}$. Our algorithm for solving (19) is called ALPCAHUS (ALPCAH for Union of Subspaces). We solve (19) via alternating minimization, and the method alternates between subspace basis estimation and data sample reassignment/clustering. We begin with subspace basis estimation.
Firstly, with data sample assignments fixed, $\forall k\in\{1,\ldots,K\}$ we estimate subspace bases by applying $T_{1}$ iterations of alternating minimization by
$\left(\bm{L}_{k}^{(T_{1})},\bm{R}_{k}^{(T_{1})},\bm{\Pi}_{k}^{(T_{1})}\right)=\operatorname*{arg\,min}_{\bm{L}_{k},\,\bm{R}_{k},\,\bm{\Pi}_{k}}$ $f_{k}(\bm{L}_{k},\bm{R}_{k},\bm{\Pi}_{k};\bm{Y}_{k})$
$\hskip 5.69054pt\text{s.t.}\hskip 5.69054pt\bm{\Pi}_{k}\succeq\alpha\bm{I}.$ (20)
The solution to (20) is described in Ref. [alpcah_journal]. However, for completeness and notational consistency, the updates are included in (21), (22), and (24). Given the current $t_{1}$ iteration, the updates for the $t_{1}+1$ iteration are given as
$\bm{L}^{(t_{1}+1)}_{k}$ $=\operatorname*{arg\,min}_{\bm{L}_{k}}f_{k}\left(\bm{L}_{k},\bm{R}^{(t_{1})}_{k},\bm{\Pi}^{(t_{1})}_{k};\bm{Y}_{k}^{(t_{2})}\right)$
$=\bm{Y}_{k}^{(t_{2})}\bm{\Pi}^{(t_{1})}_{k}\bm{R}^{(t_{1})}_{k}\left(\big(\bm{R}^{(t_{1})}_{k}\big)^{\prime}\bm{\Pi}^{(t_{1})}_{k}\bm{R}^{(t_{1})}_{k}\right)^{-1}$ (21)
$\bm{R}^{(t_{1}+1)}_{k}$ $=\operatorname*{arg\,min}_{\bm{R}_{k}}f_{k}\left(\bm{L}^{(t_{1}+1)}_{k},\bm{R}_{k},\bm{\Pi}^{(t_{1})}_{k};\bm{Y}_{k}^{(t_{2})}\right)$
$=\big(\bm{Y}_{k}^{(t_{2})}\big)^{\prime}\bm{L}^{(t_{1}+1)}_{k}\left(\big(\bm{L}_{k}^{(t_{1}+1)}\big)^{\prime}\bm{L}^{(t_{1}+1)}_{k}\right)^{-1}.$ (22)
Further, estimate the noise variance terms by
$\underbrace{\max\left(\alpha,\frac{1}{|D|}\left\|\left(\bm{Y}_{k}^{(t_{2})}-\bm{L}^{(t_{1}+1)}_{k}\big(\bm{R}_{k}^{(t_{1}+1)}\big)^{\prime}\right)\bm{e}_{i}\right\|_{2}^{2}\right)}_{p_{k}(i)}$ (23)
where $\bm{e}_{i}$ denotes the $i$th canonical basis vector that is used to select the $i$th residual column and $\alpha>0$ is a user-selected noise variance threshold parameter. Then, we update the noise variance matrix $\bm{\Pi}_{k}$ as follows
$\bm{\Pi}^{(t_{1}+1)}_{k}$ $=\operatorname*{arg\,min}_{\bm{\Pi}_{k}}f_{k}(\bm{L}^{(t_{1}+1)}_{k},\bm{R}^{(t_{1}+1)}_{k},\bm{\Pi}_{k};\bm{Y}_{k}^{(t_{2})})$
$=\text{diag}(p_{k}(1),\ldots,p_{k}(|\mathcal{C}_{k}|)).$ (24)
Because $\bm{\Pi}_{k}$ is a diagonal matrix, the $\bm{\Pi}_{k}\succeq\alpha\bm{I}$ majorization condition in (20) being equivalent to $\forall i,\lambda_{i}(\bm{\Pi}_{k})\geq\alpha$ implies that $\forall i,\nu_{i}>\alpha$. This leads to a projection to the positive set $\alpha,\infty)$ by the $\max(\alpha,\cdot)$ condition in (23). This condition is sufficient to ensure convergence as proven in [alpcah_journal, Thm. 2] and used in this work to further prove ALPCAHUS convergence in Thm. 1 using some $\alpha\in\mathbb{R}_{+}$.
Algorithm ALPCAHUS (unknown noise variances, unknown noise grouping, ensemble version with random init.)
1:Input: $\bm{Y}\in\mathbb{R}^{D\times N}$: data, $K\in\mathbb{Z}^{+}$: number of subspaces, $\{\hat{d}_{k}\in\mathbb{Z}^{+}\hskip 5.69054pt\forall k\in\{1,\ldots,K\}\}$: candidate dimension for all clusters, $q\in\mathbb{Z}^{+}$: threshold parameter, $B\in\mathbb{Z}^{+}$: base clusterings, $T_{1}\in\mathbb{Z}^{+}$: LR-ALPCAH iterations, $T_{2}\in\mathbb{Z}^{+}$: maximum alternating updates (cluster reassignments)
2:Output: $\mathcal{C}_{S}=\{c_{1},\dots,c_{N}\}$: clusters of $\bm{Y}$
3:for $b=1,\dots,B$ (in parallel) do
4:  $\mathcal{C}_{k}\sim\{1,\ldots,N\}\hskip 5.69054pt\text{s.t.}\hskip 5.69054pt|\mathcal{C}_{k}|\approx\frac{N}{\bar{K}}$ for $k=1,\dots,K$ Initialize clusters randomly
5:  $\bm{\Pi}_{k}\leftarrow\bm{I}$ for $k=1,\dots,K$ Assume homoscedastic data initially
6:  $\bm{L}_{k},\bm{R}_{k}\leftarrow\bm{U}_{:,1:\hat{d}_{k}}\bm{\Sigma}^{1/2}_{1:\hat{d}_{k},1:\hat{d}_{k}},\ \bm{\Sigma}^{1/2}_{1:\hat{d}_{k},1:\hat{d}_{k}}\bm{V}_{:,1:\hat{d}_{k}}\text{ where }\text{SVD}(\bm{Y}_{\mathcal{C}_{k}})=\bm{U}\bm{\Sigma}\bm{V}^{\prime}$ for $k=1,\dots,K$ Spectral init
7:  for $t_{2}=1,\dots,T_{2}$ (in sequence) do
8:   $\bm{L}_{k},\bm{R}_{k},\bm{\Pi}_{k}\leftarrow$ Compute (21)-(24) using $\bm{Y}_{\mathcal{C}_{k}}$ for $T_{1}$ iterations for $k=1,\dots,K$ Apply LR-ALPCAH [alpcah_journal]
9:   $\mathcal{C}_{k}\leftarrow\{\forall\bm{y}_{i}\in\bm{Y}:\forall k,$ apply (25)-(29) with $\bm{U}_{k}\text{ extracted from GramSchmidt}(\bm{L}_{k})\}$ Update cluster labels
10:   Stop execution if $\mathcal{C}_{k}^{(t_{2}+1)}=\mathcal{C}_{k}^{(t_{2})}$ is met for all clusters Stopping criteria
11:  end for
12:  $\mathcal{C}^{(b)}\leftarrow\mathcal{C}=\{c_{1},\dots,c_{N}\}$ Collect results from all trials
13:end for
@@ IV Proposed Subspace Clustering Method
[…]
Secondly, fixing the subspaces, we update the data sample assignments to clusters. The cluster update is essentially $\mathcal{C}^{\mathrm{new}}=\operatorname*{arg\,min}_{\mathcal{C}}f(\mathcal{L},\mathcal{R},\bm{\Pi},\mathcal{C}),$ with a rule to break ties in favor of the previous cluster assignment as follows. Let $\bm{U}_{k}^{(T_{1})}$ denote the Gram-Schmidt vectors of $\bm{L}_{k}^{(T_{1})}$. Define residual point error for $k$th basis by
$J_{i}(k)=\left\|\bm{y}_{i}-\bm{U}_{k}^{(T_{1})}\left(\bm{U}_{k}^{(T_{1})}\right)^{\prime}\bm{y}_{i}\right\|_{2}^{2}$ (25)
and let the set $S_{J_{i}}$ denote the set of minimizers of $J_{i}(k)$ by
$\mathcal{S}_{J_{i}}=\operatorname*{arg\,min}_{k}J_{i}(k).$ (26)
Then $\forall i$, given the label estimate $c_{i}^{(t_{2})}$ for the $t_{2}$ iteration, compute the next label by
$c_{i}^{(t_{2}+1)}=\left\{\begin{array}[]{ll}c_{i}^{(t_{2})}&\text{ if }c_{i}^{(t_{2})}\in\mathcal{S}_{J_{i}}\\ k^{*}\in\mathcal{S}_{J_{i}}&\text{ otherwise}.\end{array}\right$. (29)
@@
The solution to (29) involves the following procedure. Given a data sample $\bm{y}_{i}$, find the lowest subspace projection residual out of all subspaces by (25) and assign it to $c_{i}^{(t_{2}+1)}$ at the $t_{2}+1$ iteration. In the event of a tie, meaning there is more than one subspace that has equal residual, retain the past label $c_{i}^{(t_{2})}$ for the new label $c_{i}^{(t_{2}+1)}$. By doing this, cycling between labels is prevented for all data samples. This cluster reassignment criteria will be important to ensure convergence as proven in Thm. 1. Because (26) is with respect to $k$ and not the $i$th sample, a residual weighted with the noise parameter $\nu_{i}$ is unnecessary, since it would not change the minimizing $k$.
@@
To further improve clustering performance, consensus clustering can be leveraged over many trials. Initially, we tried using this approach with only one trial as seen in Fig. 2(b) and ALPCAHUS ($B=1$) result in Fig. 6(a). However, since $K$-subspaces in general is sensitive to initialization, during experimentation we found higher clustering accuracy by using a consensus approach with more than one trial as shown in Fig. 2(d) and ALPCAHUS ($B=16$) result in Fig. 6(a).